ACTIVITY 5.2 MEASURES OF CENTRAL TENDENCY

Calculate the Mean, Median and Mode of the following survey data. Show your solution. Follow the steps in the handouts.

1. A gardener buys 10 packets of seeds from two different companies. Each pack contains 20 seeds and he records the number of plants which grow from each pack.

COMPANY A: 20, 5, 20, 20, 20, 6, 20, 20 , 20, 8
COMPANY B: 17, 18, 15, 16, 18, 18, 17, 15, 17, 18

a. Find the mean, median, and mode for each company’s seeds.
b. Which company does the mode suggests is best?
c. Which company does the mean suggests is best?
d. Which company does the median suggests is best?

(answer and draw it into table.thankyou)
1
Expert's answer
2020-12-21T17:13:41-0500

a. Mean of company A "= \frac{\sum company \;A}{10}"

"=\frac{159}{10} \\

= 15.9"

"Median_A = \frac{(\frac{n}{2})^{th} + (\frac{n}{2}+1)^{th}}{2}

= \frac{5^{th} + (6)^{th}}{2} \\

= \frac{20 + 20}{2}

= 20"

Mode – most frequently repeated observation

"Mode_A = 20 \\

Mean_B = \frac{169}{10} \\

= 16.9 \\

Median_B = \frac{(5)^{th} + (\frac{n}{2}+1)^{th}}{2} \\

= \frac{17 + 17}{2} \\

= 17 \\

Mode_B = 18"

b. Company A mode suggest is best.

c. Company B mean suggest is best.

d. Company A median suggest is best.